OpenStudy (anonymous):

how do i find the distance between these pairs of points: (-2,-2).. (-4,-1)?

OpenStudy (anonymous):

please walk me through it

OpenStudy (anonymous):

\[\sqrt{(x-x2)(y-y2)}\]

OpenStudy (anonymous):

in english, the distance between -2 and-4 on the x axis is 2 the distance between -2 and-1 on the y axis is 1 pythatoras tells you that \[d^2=2^2+1^2=4+1=5\] so \[d=\sqrt{5}\]

OpenStudy (anonymous):

d=sqrt[(x2-x1)^2+(y2-y1)^2]

OpenStudy (anonymous):

@math pro and zbay u forgot exponents

OpenStudy (anonymous):

you can use the formula \[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\] but just make sure you 1) subtract 2) square 3) add 4) take the square root

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol sorry dont mean to be mean

OpenStudy (anonymous):

@satellite in that order? and everyone thanks for the help :)

OpenStudy (anonymous):

if u just plug the whole thing in caclulator it will work just fine too :P

OpenStudy (anonymous):

if you want to use the formula here you can write \[\sqrt{(-2-(-4))^2+(-2-(-1))^2}\] \[\sqrt{2^2+1^2}\] \[\sqrt{4+1}\] \[\sqrt{5}\]

OpenStudy (anonymous):

thanks everyone!!!!!

OpenStudy (anonymous):

first subtract in each coordinate. be careful of the signs, and make sure you are subtracting x's from x's and y's from y's

OpenStudy (anonymous):

then square each one then add then take the square root. don't try any shortcuts!

OpenStudy (anonymous):

thanks